A rain gutter is made from sheets of aluminum that are 30 inches wide. The edges are truned up to fom right angles. Determine the depth of the gutter that will allow a cross-sectional area of 40 square inches. Show that there are two different solutions to the problem. tound to the nearest tenth of an inch. The width of the gutter, after turning up the edges, is given as 30-2x. Its depth is given as x
The area is x(30-2x)
Set this expression equal to the given area, 40
The final answer is x approximately = 1.5, 13.5 inches. How did they get the final answer?

Answer by nyc_function(2741)

x(30-2x) = 40
30x - 2x^2 = 40
2x^2 - 30x + 40 = 0
At this point, use the quadratic formula to find the answers for x.
Can you take it from here?